OpenStudy (anonymous):

How do I found the EXACT value of the following? 1. cos (105 degrees) 2. sin (pi/12) cos (5 pi/12) + cos (pi/12) sin (5pi/12) 3. cos (2pi/3 - pi/6) 4. sin (5pi/12)

OpenStudy (anonymous):

use sum formulas

OpenStudy (anonymous):

cos(45+60) would work for the first one

OpenStudy (anonymous):

I don't know what sum formulas are.

OpenStudy (anonymous):

They should be in your trig book

OpenStudy (anonymous):

in that section

OpenStudy (anonymous):

I don't have a book!

OpenStudy (anonymous):

:( google sum and difference identities the idea is to make addition or subtraction the number inside the () for ones like the first you are trying to make them numbers you know on the unit circle

OpenStudy (anonymous):

Im sure if you plug those into wolfram alpha they will show you a step by step solution.

OpenStudy (callisto):

cos105 = cos(45+60) = cos45cos60-sin45sin60 Just put sin45 = cos45 = (sqrt2)/2 , cos60 =1/2 and sin60 =(sqrt3)/2 to the above step , you can get the exact value for it

OpenStudy (anonymous):

Question. Why is it 45 and 60 and not 50 and 55 or some other random number?

OpenStudy (anonymous):

because you know 45 and 60

OpenStudy (callisto):

Because they are the special angle for trigo function and its exact values are well-known

OpenStudy (anonymous):

those are special angles

OpenStudy (callisto):

* functions and their exact values

OpenStudy (anonymous):

Another question? Where are you getting (sqrt2)/2 , cos60 =1/2 and sin60 =(sqrt3)/2 to

OpenStudy (anonymous):

those are the values for those special angles

OpenStudy (anonymous):

Where can I find a list of all these values?

OpenStudy (anonymous):

you should memorize those.

OpenStudy (anonymous):

umm sec

OpenStudy (anonymous):

use triangles

OpenStudy (anonymous):

eg. unit triangle 1,1,sqrt(2) has angles 45 degrees

OpenStudy (callisto):

I can type you one here 30 45 60 sin | 1/2 | (sqrt2)/2 | (sqrt3)/2 ------------------------------------ cos | (sqrt3)/2| (sqrt2)/2 |1/2 ------------------------------------- tan | 1/sqrt3 | 1 | sqrt3

OpenStudy (callisto):

Sorry i'm always slow :S

OpenStudy (anonymous):

and half an equilateral 1,sqrt(3),2 has angles 60 and 30

OpenStudy (anonymous):

Ah I see. Let me try to get the answer and I'll tell you what I get.

OpenStudy (callisto):

Glad to hear!! Wait for your answer then :)

OpenStudy (anonymous):

 sin(45) = cos(45) = 1/sqrt(2)